OpenStudy (anonymous):

(1-tanx)/(1+tanx) = (cotx-1)/(cotx-1) only work on one side

OpenStudy (anonymous):

(1-tanx)/(1+tanx) = 1 1 - tanx = 1 + tanx tan x = 0

OpenStudy (anonymous):

typo????

OpenStudy (anonymous):

\[LHS=\frac{1-\frac{1}{cotx}}{1+\frac{1}{cotx}}\] \[\frac{\frac{cotx-1}{cotx}}{\frac{cotx+1}{cotx}}\] \[\frac{cotx-1}{\cancel{cotx}}*\frac{\cancel{cotx}}{cotx+1}\] \[\frac{cotx-1}{cotx+1}=RHS\]

OpenStudy (anonymous):

No typo

OpenStudy (anonymous):

Great! Thanks for your help!